An electron is accelerated from rest across the gap of a capacitor (two parallel plates charged -Q and +Q respectively). A hole in the top plate allows the electron to emerge with a constant velocity of v = 71 m/s. If the gap between the plates is d = 0.47 m, what is the magnitude of the electric field between the plates?
Answer #1

From

1/2 mv^2 =q delta v

delta v =(0.5*9.1*10^-31*71^2)/(1.6*10^-19)

=1.43*10^-8

E=V/d = (1.43*10^-8)/(0.47)

E=3.04*10^-8 V/m

I hope help you !!